OpenStudy (anonymous):

integration of: sqrt( (1+x)/(1-x) ) I tried using x=cos x but the answer seems to be diff.

OpenStudy (anonymous):

I got the answer as: -cos inverse (x) - sin ( cos inverse (x) ) + c which is wrong. The correct answer is: sin inverse (x) - sqrt(1-x^2) + C

hartnn (hartnn):

try x=cos 2y what you get after simplifying the square root ?

OpenStudy (anonymous):

still no use, since the first term of the answer is arcsin(x) , i think either the answer given is wrong or we have to take t=sinx , with t=sinx i dont know how to do?

OpenStudy (raden):

try integral by u-subs let u=1-x --------> x=1-u du = dx = int(1+x)/(1-x) dx = int (1+1-u)/u du = int (2-u)/u du = int 2/u du - int 1 du = ...

hartnn (hartnn):

what about sqrt(...) :P

OpenStudy (raden):

haha, misread :D

OpenStudy (anonymous):

\[\int\sqrt{\frac{1+x}{1-x}}\cdot\sqrt{\frac{1-x}{1-x}}dx\\ \int\frac{\sqrt{1-x^2}}{1-x}dx\] \[u=1-x\Rightarrow x=1-u\\ du=-dx\] \[-\int\frac{\sqrt{1-(1-u)^2}}{u}du\] \[1-u=\sin t\\ -du=\cos t\;dt\] \[\int\frac{\sqrt{1-\sin^2t}}{1-\sin t}\cos t\;dt\\ \int\frac{\cos^2t}{1-\sin t}dt\\ \int\frac{1-\sin^2t}{1-\sin t}dt\\ \int\frac{(1+\sin t)(1-\sin t)}{1-\sin t}dt\\ \int(1+\sin t)\;dt\\ t-\cos t+C\] And make sure to sub back.

OpenStudy (anonymous):

brilliant. thanks a lot